# latexText page

### v1

@@ IV-A Image Deblurring and Super-Resolution
For the image deblurring task, we first degraded the test images by convolving them with two different $9\times 9$ kernels: a uniform blur and a Gaussian blur with a standard deviation of $1.6$. In the image super-resolution task, we generated the low-resolution image by convolving the high-resolution with a $7\times 7$ Gaussian blur with a standard deviation of $1.6$, followed by downsampling with a factor of $2$. We added $5\%$ salt-and-pepper noise to the degraded images. [^2] We first randomly chose $5$% of the pixels and set them to $1$. Then, we randomly chose another $5$% of the pixels and set them to 00. The supplementary material also reports the performance of using RNP with different noise levels. We set the sketch size $K=100$ and tolerance $\epsilon=10^{-4}$ in Algorithm 1 and PCG, respectively. Similar to [12], the restoration of an RGB image was conducted by converting it to the YCbCr color-space first. Next, we applied the reconstruction algorithm to the luminance channel only and then transformed the result back to the RGB domain. For the deblurring task, we simply applied a median filter to the chroma channels for denoising. For the super-resolution task, we first used a median filter to denoise the chroma channels and then applied bicubic interpolation to upscale them to the desired resolution. Moreover, the peak signal-to-noise ratio (PSNR) values of RGB images were evaluated on the luminance channel only. We used $\ell_{p}-\ell_{1}$ to denote the method with different $p$. For the one using RNP, we added “P-”, i.e., $\bm{\mathrm{P}}-\ell_{p}-\ell_{1}$.
@@
Figure 1: The reconstructed images for different $p$ on the image deblurring and super-resolution tasks. The PSNR value is labeled at the right bottom corner of each image. The first and fifth columns are the noisy measurement and the ground truth, respectively. The first (respectively, second) row is the reconstructed starfish (respectively, leaves) image at $10$th (respectively, $6$th) iteration for the image deblurring task with uniform (respectively, Gaussian) blur. The third row is the reconstructed bike image at $7$th iteration for the image super-resolution task with a downsampling factor $2$ and $5$% salt-and-pepper noise.
TABLE II: Comparison of PSNR and ST for different $p$ in the image deblurring task with uniform and Gaussian blurs. We ran IRM $20$ iterations on each image and presented the highest PSNR values within these $20$ iterations along with the associated ST values. The highest PSNR and ST values for each image are marked in bold.
@@
TABLE III: Comparison of PSNR and ST for different $p$ in the image super-resolution task. We ran IRM $20$ iterations on each image and presented the highest PSNR values within these $20$ iterations along with the associated ST values. The highest PSNR and ST values for each image are marked in bold.
@@
IV-A1 Image Deblurring
(a) Uniform blur.
(b) Uniform blur.
(c) Gaussian blur.
(d) Gaussian blur.
(a) Starfish & uniform blur.
(b) Leaves & Gaussian blur.
Figure 2 presents the PSNR values versus iteration and wall time for different $p$. The first (respectively, second) row of Figure 2 was tested on the starfish (respectively, leaves) image with uniform (respectively, Gaussian) blur. Figures 2(a) and 2(c) show a small $p$ yielded a higher PSNR than a large $p$. It is not supervised because a small $p$ is more robust to outliers [64]. However, (3) becomes nonconvex for $p<1$, so solving (3) is more challenging. Indeed, Figures 2(b) and 2(d) illustrate that a small $p$ required much more wall time than a large one. Figure 2 illustrates that using RNP significantly accelerated the convergence speed in terms of wall time. The time for computing RNP was included in the whole wall time. From Figures 2(b) and 2(d), we even saw $\mathrm{P}-\ell_{0.5}-\ell_{1}$ converged faster than $\ell_{1}-\ell_{1}$ in terms of the wall time illustrating the effectiveness of using RNP. Figure 3 describes the number of iterations within CG at each iteration of IRM for different $p$, with the uniform and Gaussian kernels for the starfish and leaves images, respectively. Using RNP reduced the number of iterations required by CG more than $90$%, while achieving the same solution accuracy. The first and second rows of Figure 1 present the noisy, reconstructed, and ground truth images, where the degradation is reduced significantly.
@@ IV-A1 Image Deblurring
where $\mathrm{Time}_{\text{w}}$ and $\mathrm{Time}_{\text{w/o}}$ denote the wall time with and without using RNP, respectively. If $\mathrm{ST}$ is close to $1$, then using RNP significantly reduced the wall time. To study the performance of RNP further, we tested on additional $8$ images. We ran IRM $20$ iterations for different $p$. Table II presents the highest PNSR and the associated ST value for each image within $20$ iterations. Using RNP saved more than $95\%$ wall time than without RNP, illustrating the effectiveness and efficiency of using RNP. Moreover, Table II also shows that a smaller $p$ yielded a higher PSNR.
@@ IV-A2 Image Super-Resolution
Figure 5: Number of iterations within CG for different $p$ in the super-resolution task. The downsampling factor is $2$. The legend is identical to Figure 4.
Figure 4 shows the PSNR values for different $p$ in the image super-resolution of the bike image. The results are consistent with the trends in Section IV-A1, where a small $p$ yielded a higher PSNR than a larger one and using RNP significantly accelerated the convergence speed in terms of wall time. Moreover, Figure 5 presents the number of CG iterations for different $p$; the preconditioned one required fewer iterations illustrating the effectiveness of using RNP. The third row of Figure 1 presents the reconstructed images, where we observed that a small $p$ yielded a higher quality image. We tested on additional $8$ images to study the performance of using RNP further. Table III shows using RNP saved almost $70\%$ (respectively, $95\%$) time for $p=1$ (respectively, $p=0.5$) in line with the observation in Section IV-A1.
@@ IV-A3 The Choice of Sketch Size
Figure 6: Comparison of different $K$ in the image deblurring task with Gaussian kernel on the leaves image. (a): the number of iterations within CG in each iteration; (b): PSNR values versus wall time.
@@
Figure 6 presents the number of iterations within CG in the image deblurring task with Gaussian blur on the leaves image for different $K$. For $K>20$, the number of iterations within CG did not change significantly. In general, a larger $K$ would yield a more effective preconditioner, but it also increases the computational cost to compute and apply. Figure 6 shows the reconstructed PSNR values versus wall time, where clearly $K=300$ required more wall time than the others. Moreover, $K=50$ and $K=100$ performed better than the others. Since $K=100$ performed slightly better than $K=50$, we simply set $K=100$ in this subsection.
@@ IV-B Computed Tomography Reconstruction
We studied the performance of using RNP for CT reconstruction with three different regularizers: wavelet, TV with $\phi=1$, and HSϕ norm. For HSϕ, we mainly studied $\phi=1$. We used the “daub4” wavelet with $4$ levels in the “pywt” [66] toolbox. The Operator Discretization Library (ODL) [67] was used for the CT forward model. We investigated parallel-beam and fan-beam acquisition geometries with $100$ projection views. We scaled two slices from subject “067” in the AAPM CT Grand Challenge data [68] to $[0,1]$ as the test images; see Figure 7. Due to the page limit, the supplement shows the reconstruction results for Figure 7 and for the fan-beam geometry. Again we use “w/o” to denote results without using RNP, while “w-$K$” denotes RNP with sketch size $K$. Table IV summaries the wall time for computing RNP for different acquisitions with various regularizers on both CPU and GPU computational platforms, clearly highlighting the advantage of using the batch mode and the GPU computational platforms. In general, a larger $K$ would yield a faster convergence in terms of iterations than a small $K$. However, large $K$ would also increase the computation of applying RNP. Here we set $K=20$ for wavelet and TV-based reconstruction, and $K=100$ for HS1-based reconstruction, because these values worked well in practice.
@@
TABLE IV: Comparison of wall time for obtaining RNP in computed tomography reconstruction on CPU and GPU.
Parallel beam Fan beam
Wavelet TV HS Wavelet TV HS
$K$ $20$ $20$ $100$ $20$ $20$ $100$
GPU $0.11$s $0.10$s $0.40$s $0.11$s $0.09$s $0.46$s
CPU $0.48$s $0.40$s $4.20$s $0.48$s $0.52$s $4.25$s
Figure 9 presents the comparison of using RNP for wavelet based CT reconstruction. Figures 9 and 9 shows that w-$20$ converged faster than w/o in terms of iterations, illustrating the effectiveness of using RNP. Moreover, Figures 9 and 9 shows that w-$20$ converged faster than w/o in terms of wall time, demonstrating the wall time for solving the related WPM with Theorem 1 is negligible. Figure 10 presents the results of TV-based reconstruction that shows similar trends as in wavelet-based regularization. Figure 8 illustrates the reconstructed images at iterations $10$, $20$, $40$, and $60$, for both w-$K$ and w/o.
Refer to caption10101010w/o23.623.623.623.6Refer to caption27.227.227.227.2wRefer to caption20202020w/o28.228.228.228.2Refer to caption32.332.332.332.3wRefer to caption40404040w/o33.233.233.233.2Refer to captionw34.334.334.334.3Refer to caption60606060w/o34.234.234.234.2Refer to captionw34.734.734.734.7WaveletRefer to caption23.623.623.623.6Refer to caption25.925.925.925.9Refer to caption28.428.428.428.4Refer to caption31.231.231.231.2Refer to caption33.233.233.233.2Refer to caption33.633.633.633.6Refer to caption33.733.733.733.7Refer to caption33.733.733.733.7TVRefer to caption23.623.623.623.6Refer to caption29.729.729.729.7Refer to caption28.228.228.228.2Refer to caption33.933.933.933.9Refer to caption33.433.433.433.4Refer to caption36.236.236.236.2Refer to caption35.135.135.135.1Refer to caption37.037.037.037.0HS1
Figure 8: The parallel-beam CT reconstructed images with wavelet, TV, and HS1 based regularization at iterations $10$, $20$, $40$, and $60$. Columns $1$, $3$, $5$, and $7$ (respectively, $2$, $4$, $6$, and $8$) show the reconstructions without (respectively, with) RNP. The associated PSNR values are listed at the right corner of each image.
Figure 9: Comparison of using RNP for wavelet based CT reconstruction. w/o denotes the one without using RNP.
Figure 10: Comparison of using RNP for TV based CT reconstruction. w/o denotes the one without using RNP.
Figure 11: Comparison of using RNP for HS1 based CT reconstruction. w/o denotes the one without using RNP.

### v2

@@ IV-A Image Deblurring and Super-Resolution
For the image deblurring task, we first degraded the test images by convolving them with two different $9\times 9$ kernels: a uniform blur and a Gaussian blur with a standard deviation of $1.6$. In the image super-resolution task, we generated the low-resolution image by convolving the high-resolution with a $7\times 7$ Gaussian blur with a standard deviation of $1.6$, followed by downsampling with a factor of $2$. We added $5\%$ salt-and-pepper noise to the degraded images. [^2] We first randomly chose $5\%$ of the pixels and set them to $1$, and then randomly selected another $5\%$ of the pixels and set them to 0. The supplementary material also reports the performance of using RNP with different noise levels. We set the sketch size $K=100$ and tolerance $\epsilon=10^{-4}$ in Algorithm 1 and PCG, respectively. Similar to [12], the restoration of an RGB image was conducted by converting it to the YCbCr color-space first. Next, we applied the reconstruction algorithm to the luminance channel only and then transformed the result back to the RGB domain. For the deblurring task, we simply applied a median filter to the chroma channels for denoising. For the super-resolution task, we first used a median filter to denoise the chroma channels and then applied bicubic interpolation to upscale them to the desired resolution. Moreover, the peak signal-to-noise ratio (PSNR) values of RGB images were evaluated on the luminance channel only. We used $\ell_{p}-\ell_{1}$ to denote the method with different $p$. For the one using RNP, we added “P-”, i.e., $\bm{\mathrm{P}}-\ell_{p}-\ell_{1}$.
IV-A1 Image Deblurring
(a) Uniform blur.
(b) Uniform blur.
(c) Gaussian blur.
(d) Gaussian blur.
(a) Starfish & uniform blur.
(b) Leaves & Gaussian blur.
@@ IV-A1 Image Deblurring
Figure 3: The reconstructed images for different $p$ on the image deblurring (first and second rows) and super-resolution (third row) tasks. The PSNR value is labeled at the right bottom corner of each image. The first and fifth columns are the noisy measurement and the ground truth, respectively. The first (respectively, second) row is the reconstructed starfish (respectively, leaves) image at $10$th (respectively, 66th) iteration for the image deblurring task with uniform (respectively, Gaussian) blur. The third row is the reconstructed bike image at $7$th iteration for the image super-resolution task with a downsampling factor 22 and 55% salt-and-pepper noise.
TABLE III: Comparison of PSNR and ST for different $p$ in the image deblurring task with uniform and Gaussian blurs. We ran IRM $20$ iterations on each image and presented the highest PSNR values within these 2020 iterations along with the associated ST values. The highest PSNR and ST values for each image are marked in bold.
@@
TABLE IV: Comparison of PSNR and ST for different $p$ in the image super-resolution task. We ran IRM $20$ iterations on each image and presented the highest PSNR values within these 2020 iterations along with the associated ST values. The highest PSNR and ST values for each image are marked in bold.
@@
Figure 1 presents the PSNR values versus iteration and wall time for different $p$. The first (respectively, second) row of Figure 1 was tested on the starfish (respectively, leaves) image with uniform (respectively, Gaussian) blur. Figures 1(a) and 1(c) show a small $p$ yielded a higher PSNR than a large $p$. It is not surprising because a small $p$ is more robust to outliers [66]. However, (3) becomes nonconvex for $p<1$, so solving (3) is more challenging. Indeed, Figures 1(b) and 1(d) illustrate that a small $p$ required much more wall time than a large one. Figure 1 illustrates that using RNP significantly accelerated the convergence speed in terms of wall time. The time for computing RNP was included in the whole wall time. From Figures 1(b) and 1(d), we even saw $\mathrm{P}-\ell_{0.5}-\ell_{1}$ converged faster than $\ell_{1}-\ell_{1}$ in terms of the wall time illustrating the effectiveness of using RNP. Figure 2 describes the number of iterations within CG at each iteration of IRM for different $p$, with the uniform and Gaussian kernels for the starfish and leaves images, respectively. Using RNP reduced the number of iterations required by CG more than 9090%, while achieving the same solution accuracy. The first and second rows of Figure 3 present the noisy, reconstructed, and ground truth images, where the degradation is reduced significantly.
@@
where $\mathrm{Time}_{\text{w}}$ and $\mathrm{Time}_{\text{w/o}}$ denote the wall time with and without using RNP, respectively. If $\mathrm{ST}$ is close to $1$, then using RNP significantly reduced the wall time. To study the performance of RNP further, we tested on additional $8$ images. We ran IRM $20$ iterations for different $p$. Table III presents the highest PNSR and the associated ST value for each image within $20$ iterations. Using RNP saved more than $95\%$ wall time than without RNP, illustrating the effectiveness and efficiency of using RNP. Moreover, Table III also shows that a smaller $p$ yielded a higher PSNR.
@@ IV-A2 Image Super-Resolution
Figure 5: Number of iterations within CG for different $p$ in the super-resolution task. The downsampling factor is 22. The legend is identical to Figure 4.
Figure 4 shows the PSNR values for different $p$ in the image super-resolution of the bike image. The results are consistent with the trends in Section IV-A1, where a small $p$ yielded a higher PSNR than a larger one and using RNP significantly accelerated the convergence speed in terms of wall time. Moreover, Figure 5 presents the number of CG iterations for different $p$; the preconditioned one required fewer iterations illustrating the effectiveness of using RNP. The third row of Figure 3 presents the reconstructed images, where we observed that a small $p$ yielded a higher quality image. We tested on additional $8$ images to study the performance of using RNP further. Table IV shows using RNP saved almost $70\%$ (respectively, $95\%$) time for $p=1$ (respectively, $p=0.5$) in line with the observation in Section IV-A1.
@@ IV-A3 The Choice of Sketch Size
(a) $p=0.5.$
(b) $p=0.5.$
Figure 6: Comparison of different $K$ in the image deblurring task with Gaussian kernel on the leaves image for $p=0.5$. (a): the number of iterations within CG in each iteration; (b): PSNR values versus wall time.
@@
Figure 6(a) presents the number of iterations within CG in the image deblurring task with Gaussian blur and $p=0.5$ on the leaves image for different $K$. For $K>20$, the number of iterations within CG did not change significantly. Notice that at the $7$th iteration, $K=20$ requires significantly more CG iterations than other iterations, where the preconditioner is not very effective. However, the required number of CG iterations is still much lower than in the case without a preconditioner. In general, a larger $K$ would yield a more effective preconditioner, but it also increases the computational cost to compute and apply. Figure 6(b) shows the reconstructed PSNR values versus wall time, where clearly $K=300$ required more wall time than the others. Moreover, $K=50$ and $K=100$ performed better than the others. Since $K=100$ performed slightly better than $K=50$, we simply set $K=100$ in this subsection.
IV-A4 Comparison of GKS and ADMM
We compared the performance of IRM with RNP (w-$20$), GKS [30], and ADMM for $p=0.5,1$ on the super-resolution task with the bike image. Figure 7 presents the cost values in terms of number of iterations and wall time. Figures 7(a) and 7(c) show that w-$20$ is the fastest algorithm in terms of iterations. Compared to w-$20$, GKS requires significantly less computation per iteration. Indeed, Figure 7(b) indicates GKS converged fastest than the other methods in terms of wall time for $p=1$. However, for $p=0.5$, we observed that w-$20$ converged faster than GKS in terms of wall time illustrating the effectiveness of using RNP for acceleration when the problem becomes more challenging.
(a) $p=1.$
(b) $p=1.$
(c) $p=0.5.$
(d) $p=0.5.$
Figure 7: Comparison with GKS and ADMM on the image super-resolution task on the bike image with $p=0.5,1$.
@@ IV-B Computed Tomography Reconstruction
We studied the performance of using RNP for CT reconstruction with three different regularizers: wavelet, TV with $\phi=1$, and HSϕ norm. Moreover, we also compared our method with the ADMM algorithm [38]. For HSϕ, we mainly studied $\phi=1$. We used the “daub4” wavelet with $4$ levels in the “pywt” [68] toolbox. The Operator Discretization Library (ODL) [69] was used for the CT forward model. We investigated both parallel-beam and fan-beam acquisition geometries with $100$ projections. For both geometries, the reconstruction space was uniformly discretized on $20,20]\times[-20,20]\,\mathrm{cm}^{2}$ with a resolution of $512\times 512$. In the parallel-beam case, projection data were collected over uniformly spaced views in $0,180^{\circ}]$, using a linear detector with $1024$ bins spanning $40,40]\,\mathrm{cm}$. In the fan-beam case, projection data were acquired over uniformly spaced views in $0,360^{\circ}]$, with a linear detector of $1024$ bins covering $60,60]\,\mathrm{cm}$. We scaled two slices from subject “067” in the AAPM CT Grand Challenge data [70] to $[0,1]$ as the test images; see Figure 9. The projections were generated by applying the forward model and then adding Gaussian noise with zero mean and variance $10^{-4}$. Due to the page limit, the supplement shows the reconstruction results for Figure 9 and for the fan-beam geometry. Again we use “w/o” to denote results without using RNP, while “w-KK” denotes RNP with sketch size $K$.
Table II summarizes the wall time for computing RNP for different acquisitions with various regularizers on both CPU and GPU computational platforms, clearly highlighting the advantage of using the batch mode and the GPU computational platforms. In general, a larger $K$ would yield a faster convergence in terms of iterations than a small $K$. Figure 8 presents the values of $L_{f}^{\bm{\mathrm{P}}}\|\bm{\mathrm{x}}^{1}-\bm{\mathrm{x}}^{*}\|_{\bm{\mathrm{P}}}^{2}$ along the change of $K$ with fan-beam and parallel-beam acquisitions. Figure 8 shows that by using RNP, we can significantly reduce $L_{f}^{\bm{\mathrm{P}}}\|\bm{\mathrm{x}}^{1}-\bm{\mathrm{x}}^{*}\|_{\bm{\mathrm{P}}}^{2}$ that one can expect an acceleration. In this case, we observe that RNP becomes less effective at reducing $L_{f}^{\bm{\mathrm{P}}}\|\bm{\mathrm{x}}^{1}-\bm{\mathrm{x}}^{*}\|_{\bm{\mathrm{P}}}^{2}$ when $K>100$. Moreover, a larger $K$ may slightly increase $L_{f}^{\bm{\mathrm{P}}}\|\bm{\mathrm{x}}^{1}-\bm{\mathrm{x}}^{*}\|_{\bm{\mathrm{P}}}^{2}$ but still remains effective at reducing it. In practice, large $K$ would also increase the computation of applying RNP. Here we set $K=20$ for wavelet and TV-based reconstruction, and $K=100$ for HS1-based reconstruction, because these values worked well in practice.
Figure 8: $L_{f}^{\bm{\mathrm{P}}}\|\bm{\mathrm{x}}^{1}-\bm{\mathrm{x}}^{*}\|_{\bm{\mathrm{P}}}^{2}$ values versus different $K$ for fan-beam and parallel-beam acquisitions.
@@
Figure 11 presents the comparison of using RNP for wavelet based CT reconstruction. Figures 11 and 11 shows that w-$20$ converged faster than w/o in terms of iterations, illustrating the effectiveness of using RNP. Moreover, Figures 11 and 11 shows that w-$20$ converged faster than w/o in terms of wall time, demonstrating the wall time for solving the related WPM with Theorem 1 is negligible. Figure 12 presents the results of TV-based reconstruction that shows similar trends as in wavelet-based regularization. Obviously, ADMM is the slowest algorithm in these comparisons. Figure 10 illustrates the reconstructed images at iterations $10$, 2020, $40$, and $60$, for both w-$K$ and w/o.
Refer to caption1010w/o23.623.6Refer to caption27.227.2wRefer to caption2020w/o28.228.2Refer to caption32.332.3wRefer to caption4040w/o33.233.2Refer to captionw34.334.3Refer to caption6060w/o34.234.2Refer to captionw34.734.7WaveletRefer to caption23.623.6Refer to caption25.925.9Refer to caption28.428.4Refer to caption31.231.2Refer to caption33.233.2Refer to caption33.633.6Refer to caption33.733.7Refer to caption33.733.7TVRefer to caption23.623.6Refer to caption29.729.7Refer to caption28.228.2Refer to caption33.933.9Refer to caption33.433.4Refer to caption36.236.2Refer to caption35.135.1Refer to caption37.037.0HS1
Figure 10: The parallel-beam CT reconstructed images with wavelet, TV, and HS1 based regularization at iterations $10$, 2020, $40$, and $60$. Columns 11, 33, 55, and $7$ (respectively, 22, 44, 66, and 88) show the reconstructions without (respectively, with) RNP. The associated PSNR values are listed at the right corner of each image.
Figure 11: Comparison of using RNP and ADMM for wavelet based CT reconstruction. w/o denotes the one without RNP.
Figure 12: Comparison of using RNP and ADMM for TV based CT reconstruction. w/o denotes the one without RNP.
Figure 13: Comparison of using RNP and ADMM for HS1 based CT reconstruction. w/o denotes reconstruction without RNP.
@@
IV-B1 Comparison of IRM and GKS
We compared the performance of WAPG using RNP with IRM and GKS for wavelet-based parallel-beam CT reconstruction. Figure 14 presents the cost values versus the number of iterations and wall time for different methods. For IRM and GKS, the $\ell_{1}$ norm is smoothed, but the cost values presented here are still computed with the $\ell_{1}$ norm. We observed that IRM is the fastest algorithm in terms of the number of iterations. However, since IRM requires higher computation per iteration, we found that w-$20$ is the fastest algorithm in terms of wall time. Compared to IRM, GKS significantly reduces the computational cost per iteration, which is also validated in our experiment. However, we observed that w-$20$ is still faster than GKS in terms of wall time in our setting.
Figure 14: Comparison of using RNP, IRM, and GKS for wavelet based CT reconstruction.
IV-C Computed Tomography Reconstruction with Real Data
We used the publicly available walnut CT dataset provided by the Finnish Inverse Problems Society, which was acquired with a fan-beam geometry [71]. The dataset includes a given matrix $\bm{\mathrm{A}}$ and the corresponding $120$ projection measurements $\bm{\mathrm{y}}$. We evaluated the effectiveness of RNP for image reconstruction by solving (4) with TV-based regularization, with the sketch size $K=100$. In this experiment, we subsampled the data and used only $40$ projections. Figure 15 shows the cost values versus the number of iterations and wall time, demonstrating our approach converged faster than others. Figure 16 presents the reconstructed images. Using RNP recovered sharper details than APM without RNP at the same number of iterations, illustrating the effectiveness of RNP. Because walnut data did not include the ground truth image, PSNR values are not reported.
Figure 15: Comparison of using RNP and ADMM for TV based reconstruction on the walnut data. w/o denotes reconstruction without RNP.
Refer to caption1010w/oRefer to captionwRefer to caption2020w/oRefer to captionwRefer to caption4040w/oRefer to captionwRefer to caption6060w/oRefer to captionw
Figure 16: CT reconstructions of the walnut data using TV based regularization are shown at iterations $10$, 2020, $40$, and $60$. Columns 11, 33, 55, and $7$ (respectively, 22, 44, 66, and 88) show the reconstructions without (respectively, with) RNP.